Bob is driving along a straight and level road straight toward a mountain. At some point on his trip he measures the angle of elevation to the top of the mountain and finds it to be 24°. He then drives 1 mile ( 1 mile = 5280 ft ) more and measures the angle of elevation to be 33°. Find the height of the mountain to the nearest foot.
Answer by mananth(16949)


.
0.65x= 0.45x + 2376
0.2x= 2376
x=11880
11880+5280=17160
tan 24= h/17160
h= 0.45*17160
h=7640 ft
( check calc)